What is the frequency of a photon if the energy is 7.27 x 10^-19 J? (h=6.626 x 10^-34 J x s)
1
Expert's answer
2020-04-15T01:51:28-0400

"E=h\u03bd"


"ν=\frac{E}{v}=\frac{7.27\times10^{-19} J}{6.626\times10^{-34} J\times s}= 1.00\times10^{15} s^{-1}"